OpenStudy (anonymous):

Find the volume of the solid which is bounded by the cylinder x2+y2 =4, below by the xy-plane and above by the plane z = y.

OpenStudy (anonymous):

\[\int\limits_{0}^{2\pi}\int\limits_{0}^{2}\int\limits_{0}^{rsin(\theta)} rdzdrd{\theta}\]